# Homomorphism, substructure and ideal: Elementary but rigorous aspects of renormalization group or hierarchical structure of topological orders

### v2

@@ Abstract
We propose a general quantum Hamiltonian formalism of a renormalization group (RG) flow with an emphasis on generalized symmetry by interpreting the elementary relationship between homomorphism, quotient ring, and projection. In our formalism, the noninvertible nature of the ideal of a fusion ring realizing the generalized symmetry of an ultraviolet (UV) theory plays a fundamental role in determining condensation rules between anyons, resulting in the infrared (IR) theories. Our algebraic method applies to the domain wall problem in $2+1$ dimensional topologically ordered systems and the corresponding classification of $1+1$ dimensional gapped phase, for example. An ideal decomposition of a fusion ring provides a straightforward but strong constraint on the gapped phase with noninvertible symmetry and its symmetry-breaking (or emergent symmetry) patterns. Moreover, even in several specific homomorphisms connected under massless RG flows, less familiar homomorphisms appear, and we conjecture that they correspond to partially solvable models in recent literature. Our work demonstrates the fundamental significance of the abstract algebraic structure, ideal, for the RG in physics.
@@ I.1 Summary of the main ideas
  • We provide a rigorous interpretation of sandwich construction in literature as implications of the massless RG to the massive RG. (Fig. 3 and Eq. (9))
  • Only by assuming the UV and IR fusion rings, the homomorphism is not unique.
@@ II Substructure from homomorphism: Projection as renormalization group flow
where $\mathbf{A}$ is an algebra describing generalized symmetries (or integral of motions) or underlying fusion category of anyons, and the lower index $(1)$ and $(2)$ label the ultra-violet (UV) and infra-red (IR) theories, respectively. The corresponding argument has appeared evidently in [32] based on the earlier work [95] and the folding trick [11]. When interpreting this homomorphism as a defect between two theories, it is called the RG domain wall, and it is expected to describe a massless RG[86]. We also note that the categorical structure corresponding to the ring homomorphism, called a functor, has been studied as gapped (or symmetry-preserving) domain walls in [22, 23, 25, 27] with emphasis on the theoretical aspect of TQFTs or TOs.
@@
We restrict our attention to abstract algebra, not to category theory, because the underlying category theories can be premodular fusion or superfusion categories, or even outside of them that are still under development. However, the basic fusion ring in this manuscript is well-defined and has been studied several times both in the mathematics and physics communities. In other words, our discussion can be a clue to constructing a new fundamental class of categories. The fusion rings with negative or fractional coefficients are the corresponding structure, and they appeared in the study of fermionic string theories [119] and more recently in [88, 57], for example. The corresponding fusion rings have been studied more generally by the first author in [89, 90] with a close connection to the simple current extension by Schellekens and the collaborators[72, 73, 74, 75, 76, 77, 78, 79]. We also note that the topological symmetry operator (or integral of motion) is an operator over complex fields. Hence, the quantum Hamiltonian formalism of the symmetry operator itself does not restrict the coefficients to nonnegative integer coefficients of the fusion rings. The appearance of a nonnegative integer coefficient is the consequence of the correspondence between the symmetry operator and the defects under the topological Wick rotation (or modular $S$ transformation in two-dimensional CFTs)[131] [^1] We thank Yuji Tachikawa, Jurgen Fuchs, and Shinichiro Yahagi for the related discussions..
@@
A gapped phase is an ideal of unbroken subring A’. (8)
@@ II.1 Unbroken and emergent symmetry
Based on the arguments in this section, we provide a definition of unbroken (or exact) symmetry and emergent symmetry under an RG flow $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$. A phenomenology of unbroken symmetry under RG flow is simple: a symmetry subring $\mathbf{A}_{(1)}^{\text{ub}}\subset\mathbf{A}_{(1)}$ of UV theory, which does not change under a RG flow $\rho$, is an unbroken symmetry. In the algebraic language, this condition means that $\mathbf{A}_{(1)}^{\text{ub}}$ and $\text{Ker}\rho$ are disjoint except for 0. Hence, the following defines the unbroken symmetry,
@@
where dim is the dimension of a ring in linear algebra. By taking a maximal unbroken subalgebra under $\rho$, one can restrict the form of perturbation in the UV theory, and the above inequality provides the constraint in comparing the IR (or emergent) theory.
@@
For the readers interested in higher-dimensional systems, we note that the above arguments themselves can be applied straightforwardly to $D$-dimensional systems when treating the generalized symmetries as integral of motion or symmetry operators. When the IR theory becomes an extended TQFT under massive perturbations, the subring structure formed by $d$-dimensional symmetry operators in UV theory, where d=0,,Dd=0,...,D specifies dimensionality of symmetry. This ring structure will be obtained from the application of the topological Wick rotation to composite defects in [188], and the subring symmetry of them will appear as symmetry of the IR theory. If the corresponding flows are massless, the quotient ring of UV theory appears as IR symmetry.
@@ III Generating ideal: Role of nonabelian anyon or noninvertible symmetry
One can construct possible IR theories only from the algebraic data of a UV theory, but the construction of the homomorphisms between given UV and IR theories requires more involved calculation and consistency checks. In the next section, we solve this problem in several models.
@@ IV Construction of homomorphism
Before moving into the respective discussions, we remark on one of the most nontrivial points of our findings. Surprisingly, even in the flow from the tricritical Ising model to the Ising model, the homomorphism is not unique. However, by requiring a variant of integrability, the solution becomes unique. This phenomenon itself is analogous to ideal number decomposition by Kummer, and the integrability will play a fundamental role in constructing a treatable class of models. Inversely, the other homomorphisms are mysterious to some extent, but there exists a conjectural phenomenology to them. If one assumes the IR theory $\mathbf{A}_{(2)}$ or consequence of RG flow is integrable, one can say this homomorphism induces partially integrable embedding of the theory to UV theory $\mathbf{A}_{(1)}$, $\mathbf{A}_{(2)}\subset\mathbf{A}_{(1)}$. Fortunately, there exist very similar phenomena called partially solvable systems in recent literature[91, 92, 93] [^3] We thank Chihiro Matsui for introducing us to the idea of this embedding and related ideas in the literature and Hosho Katsura for indicating useful references in this research field.. Historical aspects of those models have been summarized in the introduction of [93], and we note the Hilbert space fragmentation as a fundamental aspect, and note a recent review [239]. We conjecture that a ring homomorphism generally corresponds to integrable or partially solvable systems. Further analysis of the RG flows by perturbative QFT methods is an interesting direction, and it will be described by UV CFT with multiple perturbations as in [240].
@@ IV.3 $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$
Here we study the homomorphism induced from the flow $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$. We use the same notation in the previous subsection, and we distinguish the simple currents of each $SU(2)_{1}$ by writing them as $J^{[i]}$, $i=1,2,3$. In $SU(2)_{3}$ there are four primary fields with spins ranging from 0 to $\frac{3}{2}$, which we list in ascending order of spin as $I^{\prime},x^{\prime},\tau^{\prime},J\textquoteright$}. Here, $\tau^{\prime}$ is the primary field that satisfies the Fibonacci fusion rule $\tau^{\prime}\times\tau^{\prime}=1+\tau^{\prime}$, and $J\textquoteright$ is the simple current of $SU(2)_{3}$. Here we study the homomorphism preserving $Z_{2}$ structures $\{I,J^{[1]}J^{[2]}J^{[3]}\}$ and $\{I,J^{\prime}\}$.
@@ 11 $0+1+2$ $(1/2)+(3/2)$ 11
$\phi_{2}(=J\textquoteright)$ is the $\mathbb{Z}_{2}$ simple current of the theory. By computing the $Z_{2}$ charge of each operator concerning $J$ or $J\textquoteright$, we can restrict the operators that should be connected by the homomorphism. In the present case, $\phi_{1/2}$ and $\phi_{3/2}$ belong to the $Z_{2}$ uncharged sector and the others belong to the $Z_{2}$ charged sector. One must note, as in the Majorana-fermion case, that $Z_{2}$ invariant operators split after the $Z_{2}$ simple current extension. In the present case the $\mathbb{Z}_{2}$-invariant sector is $\phi^{\prime}_{1}$, which therefore splits into $\phi^{\prime}_{1,0}$ and $\phi^{\prime}_{1,1}$ where the lower index 0, 11 corresponds to $Z_{2}$ fermion parity. Even after the $Z_{2}$ extension, the number of operators in the uncharged and charged sectors is significantly different, and this is in sharp contrast to the Majorana fermion case with the fermionic T-duality (or a variant of supersymmetry in older literature[245]). Moreover, in this case, the $Z_{2}$ invariant sector is uncharged.
@@
$\Psi^{\prime}_{1,1}$ $=\frac{\Phi^{\prime}_{1,0}}{\sqrt{2}}$ (55)
@@
We replace the $\Psi^{\prime}$ with the $\phi^{\prime}$ because of the ring isomorphism (or the CFT/TQFT correspondence or CFT/SymTFT correspondence). By evaluating the charge of each operator with respect to $J\textquoteright$, we find that a product of two $SU(2)_{1}$ simple currents such as $J^{\prime[i]}J^{\prime[j]}$ is mapped to a linear combination of $\phi^{\prime}_{0},\ \phi^{\prime}_{1,0},\ \phi^{\prime}_{1,1},\ \phi^{\prime}_{2}$. Thus we set
@@
In this work, we studied the explicit construction of a ring homomorphism between fusion rings that determines the possible IR symmetry systematically. Even in the well-established models connected by massless RG flows, a nontrivial structure appears. To obtain the homomorphism more generally, systematic studies on the ideal structure of the model are necessary, as we have stressed in the section. II. When the ideal is commutative, the existence of an eigenvalue 0 is guaranteed, and its existence in lattice models seems promising, in principle (However, we note a more involved problem of emanant symmetry in [141, 142, 143]). In other words, if the ideal is not commutative, this implies difficulty in obtaining the IR theory by projection, and the IR symmetry will become more difficult to realize in a lattice model without taking some limit. Further studies of the ideal structure will be fundamentally important for the classification and construction of general TOs. In this work, we have restricted our attention to fusion rings appearing in 22 22-dimensional CFT or $2+1$ dimensional TQFT, but the same construction of RG by ideal should be applicable to higher-dimensional systems. From this view, nontrivial transformations of UV and IR symmetry, as in [253, 254], are worth further study.
@@
We also note a related research direction toward coupled models with extended symmetries. By using the folding trick[11], it is known that the construction of extended model $\mathbf{F}_{\text{UV}\times\text{IR}}$ with $\mathbf{A}_{(1)}\otimes\mathbf{A}_{(2)}\subset\mathbf{F}_{\text{UV}\times\text{IR}}$ plays a fundamental role also for the construction of homomorphism[32]. In a $Z_{N}$ symmetric case, the extension is constructed by studying a common subgroup of $\mathbf{A}_{(1)}$ and $\mathbf{A}_{(2)}$. The anomaly-free object in these theories is the integer spin simple current[72, 73, 74, 75, 76, 77, 78, 79] and its nonchiral generalizations[85, 68] and the existence of this structure enables one to obtain the extended theory $\mathbf{F}_{\text{UV}\times\text{IR}}$. Hence, one expects the same strategy for the common subalgebra structure, but the gauging or extension for the product of this common subalgebra has not been studied as far as we know. As a natural extension of the $Z_{N}$ case, the nonsimple current in the coupled model forms a nonsimple current with (half) integer spin after taking the Deligne product (This condition corresponds to single-validness of the wavefunctions in TOs. The term nonsimple current appeared in the study of heterotic string theories. This indicates nontrivial extension or insertion of duality defects as in[131, 255, 256]). For this purpose, recent progress on the orbifolding procedure for the noninvertible symmetry [102, 103] might be useful. However, we stress that it is necessary to invent the corresponding extension, which is the inverse operation of the orbifolding. After the extension, the representation of the system will become intrinsically anyonic[257, 186].
@@
Finally, we make a comment on the interpretation of projection in the manuscript based on more recent literature on open quantum systems. In this manuscript, we have restricted our attention to static quantum phases and their transition, but by interpreting this projection as a measurement, our construction of the projection will produce a kind of treatable measurement-induced quantum phase transition[213, 214, 215] satisfying $c$ and $g$-theorem[258, 259, 260]. Further studies on the interpretation of the ideal structure in a quantum many-body system and its realization in an experimental setting are interesting research directions. Moreover, by interpreting the ring with a negative or fractional coefficient as a category, premodular fusion categories will appear, and they are expected to describe TOs in mixed states[261, 262, 263].

### v3

@@ Abstract
We propose a general quantum Hamiltonian formalism of a renormalization group (RG) flow with an emphasis on generalized symmetry by interpreting the elementary relationship between homomorphism, quotient ring, and projection. In our formalism, the noninvertible nature of the ideal of a fusion ring realizing the generalized symmetry of an ultraviolet (UV) theory plays a fundamental role in determining condensation rules between anyons, resulting in the infrared (IR) theories. Our algebraic method applies to the domain wall problem in $2+1$ dimensional topologically ordered systems and the corresponding classification of $1+1$ dimensional gapped phase, for example. An ideal decomposition of a fusion ring provides a straightforward but strong constraint on the gapped phase with noninvertible symmetry and its symmetry-breaking (or emergent symmetry) patterns. Moreover, even in several specific homomorphisms connected under massless RG flows, less familiar homomorphisms appear, and we conjecture that they correspond to partially solvable models in recent literature. Our work demonstrates the fundamental significance of the abstract algebraic structure (beyond group structure), ideal, for the RG in physics.
@@
In the previous century, group theory and the corresponding symmetry have been studied widely in physics communities (See [94, 95], for example). The extension of this framework to the ring theory is known as generalized symmetry[96, 97]. Our approach establishes this research direction in the quantum Hamiltonian formalism by formulating the RGs with mathematically rigorous relations. We stress that the ideal and its generalizations play the most fundamental role in studying the condensations resulting in the appearance of emergent symmetries, and this is completely beyond group theory.
@@ I.1 Summary of the main ideas
  • We provide a rigorous interpretation of the sandwich construction in literature as implications of the massless RG to the massive RG. (Fig. 3 and Eq. (11))
  • A rigorous definition of the unbroken and emergent symmetries and their dual relations has been shown (Eq.(12) and Eq. (14)).
  • Only by assuming the UV and IR fusion rings, the homomorphism is not unique (Section IV).
@@ II Substructure from homomorphism: Projection as renormalization group flow
$\rho(a_{(1)}+b_{(1)})$ $=\rho(a_{(1)})+\rho(b_{(1)}),$ (2)
$\rho(a_{(1)}\times b_{(1)})$ $=\rho(a_{(1)})\times\rho(b_{(1)}),$ (3)
where $\mathbf{A}$ is an algebra describing generalized symmetries (or integral of motions) or underlying fusion category of anyons, the lower index $(1)$ and $(2)$ label the ultra-violet (UV) and infra-red (IR) theories, respectively, and $a,b$ are the algebraic objects. The above conditions represent that the structure of fusion is compatible or preserved under the application of domain walls, and the theory is transformed smoothly (or masslessly) into another theory. The corresponding argument has appeared evidently in [32] based on the earlier work [99] and the folding trick [11]. When interpreting this homomorphism as a defect between two theories, it is referred to as the RG domain wall, and it is expected to describe a massless RG[86]. We also note that the categorical structure corresponding to the ring homomorphism, called a functor, has been studied as gapped (or symmetry-preserving) domain walls in [22, 23, 25, 27] with emphasis on the theoretical aspect of TQFTs or TOs. It should also be stressed that the underlying category theory should be generalized to $\mathbb{C}$-linear category, where $\mathbb{C}$ represents the complex number field, because the gapped domain wall can produce noninteger coefficients[87]. For readers interested in the underlying phenomenology of this phenomenon, we explain it in Appendix B and note the related work by the first author[100].
@@
We restrict our attention to abstract algebra, not to category theory, because the underlying category theories can be premodular fusion or superfusion categories, or even outside of them that are still under development. However, the basic fusion ring in this manuscript is well-defined and has been studied several times both in the mathematics and physics communities. In other words, our discussion can be a clue to constructing a new fundamental class of categories. The fusion rings with negative or fractional coefficients are the corresponding structure, and they appeared in the study of fermionic string theories [124] and more recently in [88, 57], for example. The corresponding fusion rings have been studied more generally by the first author in [89, 90] with a close connection to the simple current extension by Schellekens and the collaborators[72, 73, 74, 75, 76, 77, 78, 79].
We also note that the topological symmetry operator (or integral of motion) is an operator over $\mathbb{C}$[125]. Hence, the quantum Hamiltonian formalism of the symmetry operator itself does not restrict the coefficients to nonnegative integer coefficients, which are only permitted in the fusion category. The appearance of a nonnegative integer coefficient is the consequence of the correspondence between the symmetry operator and the defects under the topological Wick rotation (or modular $S$ transformation in two-dimensional CFTs)[125] [^1] We thank Yuji Tachikawa, Jurgen Fuchs, and Shinichiro Yahagi for the related discussions.. In other words, the fusion categories studied in (pure) mathematics have become physical research interests, by identifying the algebraic objects as defect or symmetry operator (or conserved charges) in the pioneering work [125]. Moreover, by focusing on their realizations as symmetry operators, which form rings over the complex number field $\mathbb{C}$, the ideas have been generalized to $\mathbb{C}$-linear categories. Readers who are familiar with particular types of fusion categories in physics may not be used to treating ring over $\mathbb{C}$, but this kind of structure, conserved charges in quantum Hamiltonian systems described by linear algebra, is more established in other sciences, including experimental and theoretical physics, and mathematics. Keeping in mind this fact, it is remarkable that the formalism in [125] is based on the quantum Hamiltonian formalism. We also comment on the realizations as defects, which do not permit noninteger coefficients before algebraic objects, but they are not in the main scope of this paper. The distinctions between the defect and symmetry operators and the related problems have been studied in [127, 128, 129, 130].
@@
A gapped phase is an ideal of unbroken subring $\mathbf{A^{\prime}}$. (10)
@@ II.1 Unbroken and emergent symmetry
Based on the arguments in this section, we define unbroken (or exact) symmetry and emergent symmetry under an RG flow $\rho:\mathbf{A}_{(1)}\rightarrow\mathbf{A}_{(2)}$. The discussions in this subsection provide the unified definition distinguishing the unbroken and emergent symmetry with a complete and rigorous language. The emergent symmetry and the corresponding phenomena are the central interest of both condensed matter and high-energy physics[133], and our expression will apply to a large class of models when their symmetries are described by a ring.
A phenomenology of unbroken symmetry under RG flow is simple: a symmetry subring $\mathbf{A}_{(1)}^{\text{ub}}\subset\mathbf{A}_{(1)}$ of UV theory, which does not change under a RG flow $\rho$, is an unbroken symmetry. In the algebraic language, this condition means that $\mathbf{A}_{(1)}^{\text{ub}}$ and $\text{Ker}\rho$ are disjoint except for 0. Hence, the following defines the unbroken symmetry,
@@
where dim is the dimension of a ring in linear algebra. In other words, the dimension of the emergent symmetry $\mathbf{A}_{(2)}^{\text{em}}$ can be expressed as,
$\text{dim}\mathbf{A}^{\text{em}}_{(2)}\leq\text{dim}\mathbf{A}_{(2)}-\text{dim}\mathbf{A}_{(1)}^{\text{ub}},$ (14)
By taking a maximal unbroken subalgebra under $\rho$, one can restrict the form of perturbation in the UV theory, and the above inequality provides the constraint in comparing the IR (or emergent) theories. When taking $\mathbf{A}_{(1)}^{\text{ub}}$ a maximal unbroken subalgebra, the inequality will become equality in some cases. The resultant relation in these cases is,
$\text{dim}\mathbf{A}^{\text{em}}_{(2)}=\text{dim}\mathbf{A}_{(2)}-\text{max}\left(\text{dim}\mathbf{A}_{(1)}^{\text{ub}}\right).$ (15)
In other words, the above relation defines unbroken and emergent symmetries as duals of each other in some cases. The exact settings realizing the above equation are worth further study.
@@
For the readers interested in higher-dimensional systems, we note that the above arguments themselves can be applied straightforwardly to $D$-dimensional systems when treating the generalized symmetries as integral of motion or symmetry operators. When the IR theory becomes an extended TQFT under massive perturbations, the subring structure formed by $d$-dimensional symmetry operators in a UV theory will flow to that of the IR theory, where d=0,,Dd=0,...,D specifies dimensionality of symmetry. This ring structure will be obtained from the application of the topological Wick rotation to composite defects in [193]. If the corresponding flows are massless, the quotient ring of UV theory appears as IR symmetry.
@@ III Generating ideal: Role of nonabelian anyon or noninvertible symmetry
One can construct possible IR theories only from the algebraic data of a UV theory. However, the construction of the homomorphisms between given UV and IR theories requires more involved calculations and consistency checks. In the next section, we solve this problem in several models.
III.1 Emergent defects: intrinsically emergent objects
In this subsection, we comment on the nontriviality of the defects appearing in the IR theory. As we have already demonstrated, the IR symmetry operators have been determined completely by the UV theory, under suitable assumptions. We have mainly discussed the transformation law of symmetry operators or anyons in higher dimensions and the coefficients before the objects can be non-integer by definition.
However, when studying the corresponding transformations of the defects, such an appearance of the non-integer coefficient is strictly prohibited. Hence, there can exist IR defects that cannot be related to the UV defects straightforwardly, by assuming the compatibility with the defect and bulk RG flows. Hence, we propose emergent defect (or emergent generalized symmetry in category theory) as follows,
[…]
We note that the studies on the bulk and defect RG flow have still been limited. We note several related works on the bulk and boundary (or defect) RG flow [239, 240, 241, 242, 243]. Moreover, the anomaly inflow mechanism[20] complicates the systematic studies because one cannot apply the $g$-theorem straightforwardly to the bulk and defect RG process[244].
@@ IV Construction of homomorphism
Before moving into the respective discussions, we remark on one of the most nontrivial points of our findings. Surprisingly, even in the flow from the tricritical Ising model to the Ising model, the homomorphism is not unique. However, by requiring a variant of integrability, the solution becomes unique. This phenomenon itself is analogous to ideal number decomposition by Kummer, and the integrability will play a fundamental role in constructing a treatable class of models. Inversely, the other homomorphisms are mysterious to some extent, but there exists a conjectural phenomenology to them. If one assumes the IR theory $\mathbf{A}_{(2)}$ or consequence of RG flow is integrable, one can say this homomorphism induces partially integrable embedding of the theory to UV theory $\mathbf{A}_{(1)}$, $\mathbf{A}_{(2)}\subset\mathbf{A}_{(1)}$. Fortunately, there exist very similar phenomena called partially solvable systems in recent literature[91, 92, 93] [^3] We thank Chihiro Matsui for introducing us to the idea of this embedding and related ideas in the literature and Hosho Katsura for indicating useful references in this research field.. Historical aspects of those models have been summarized in the introduction of [93], and we note the Hilbert space fragmentation as a fundamental aspect, and note a recent review [249].
We conjecture that a ring homomorphism generally corresponds to integrable or partially solvable systems. Further analysis of the RG flows by perturbative QFT methods is an interesting direction, and it will be described by UV CFT with multiple perturbations as in [250]. We stress that such research direction is still in progress because of the difficulty of treating multiple perturbations straightforwardly (See the discussions in [251, 252]), and our formalism will provide useful predictions as an alternative approach based on the quantum Hamiltonian.
@@ IV.3 $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$
Here we study the homomorphism induced from the flow $\{SU(2)_{1}\}^{3}\to SU(2)_{3}$. We use the same notation in the previous subsection, and we distinguish the simple currents of each $SU(2)_{1}$ by writing them as $J^{[i]}$, $i=1,2,3$. In $SU(2)_{3}$ there are four primary fields with spins ranging from 0 to $\frac{3}{2}$, which we list in ascending order of spin as $I^{\prime},x^{\prime},\tau^{\prime},J$’}. Here, $\tau^{\prime}$ is the primary field that satisfies the Fibonacci fusion rule $\tau^{\prime}\times\tau^{\prime}=1+\tau^{\prime}$, and $J$’ is the simple current of $SU(2)_{3}$. Here we study the homomorphism preserving $Z_{2}$ structures $\{I,J^{[1]}J^{[2]}J^{[3]}\}$ and $\{I,J^{\prime}\}$.
@@ 11 $0+1+2$ $(1/2)+(3/2)$ 11
[…]
@@
$\Psi^{\prime}_{1,1}$ $=\frac{\Phi^{\prime}_{1,1}}{\sqrt{2}}$ (60)
@@
We replace the $\Psi^{\prime}$ with the $\phi^{\prime}$ because of the ring isomorphism (or the CFT/TQFT correspondence or CFT/SymTFT correspondence). By evaluating the charge of each operator with respect to $J$’, we find that a product of two $SU(2)_{1}$ simple currents such as $J^{\prime[i]}J^{\prime[j]}$ is mapped to a linear combination of $\phi^{\prime}_{0},\ \phi^{\prime}_{1,0},\ \phi^{\prime}_{1,1},\ \phi^{\prime}_{2}$. Thus we set
@@
In this work, we established the explicit construction of a ring homomorphism between fusion rings that determines the possible IR symmetry systematically. Moreover, their implications for massive RG, or abstract algebraic representation of the sandwich constructions in recent literature, have been proposed and shown. We conclude that the fundamental structure of symmetry in physics, the group, has been generalized to a ring in the context of the RGs.
@@
We note again that, even in the well-established models connected by massless RG flows, a nontrivial structure appears. To obtain the homomorphism more generally, systematic studies on the ideal structure of the model are necessary, as we have stressed in the section. II. When the ideal is commutative, the existence of an eigenvalue 0 is guaranteed, and its existence in lattice models seems promising, in principle (However, we note a more involved problem of emanant symmetry in [148, 129, 130]). In other words, if the ideal is not commutative, this implies difficulty in obtaining the IR theory by projection, and the IR symmetry will become more difficult to realize in a lattice model without taking some limit. Further studies of the ideal structure will be fundamentally important for the classification and construction of general TOs. In this work, we have restricted our attention to fusion rings appearing in 22 22-dimensional CFT or $2+1$ dimensional TQFT, but the same construction of RG by ideal should apply to higher-dimensional systems. From this view, nontrivial transformations of UV and IR symmetry, as in [265, 266], are worth further study.
@@
We also note a related research direction toward coupled models with extended symmetries. By using the folding trick[11], it is known that the construction of extended model $\mathbf{F}_{\text{UV}\times\text{IR}}$ with $\mathbf{A}_{(1)}\otimes\mathbf{A}_{(2)}\subset\mathbf{F}_{\text{UV}\times\text{IR}}$ plays a fundamental role also for the construction of homomorphism[32]. In a $Z_{N}$ symmetric case, the extension is constructed by studying a common subgroup of $\mathbf{A}_{(1)}$ and $\mathbf{A}_{(2)}$. The anomaly-free object in these theories is the integer spin simple current[72, 73, 74, 75, 76, 77, 78, 79] and its nonchiral generalizations[85, 68] and the existence of this structure enables one to obtain the extended theory $\mathbf{F}_{\text{UV}\times\text{IR}}$. Hence, one expects the same strategy for the common subalgebra structure, but the gauging or extension for the product of this common subalgebra has not been studied as far as we know. As a natural extension of the $Z_{N}$ case, the nonsimple current in the coupled model forms a nonsimple current with (half) integer spin after taking the Deligne product (This condition corresponds to single-valuedness of the wavefunctions in TOs. The term nonsimple current appeared in the study of heterotic string theories. This indicates nontrivial extension or insertion of duality defects as in[125, 267, 268]). For this purpose, recent progress on the orbifolding procedure for the noninvertible symmetry [107, 108] may be useful. However, we stress that it is necessary to invent the corresponding extension, which is the inverse operation of the orbifolding. After the extension, the representation of the system will become intrinsically anyonic[269, 191].
@@
Finally, we comment on the interpretation of projection in the manuscript based on more recent literature on open quantum systems. In this manuscript, we have restricted our attention to static quantum phases and their transition, but by interpreting this projection as a measurement, our construction of the projection will produce a kind of treatable measurement-induced quantum phase transition[217, 218, 219] satisfying $c$ and $g$-theorem[270, 271, 272]. Further studies on the interpretation of the ideal structure in a quantum many-body system and its realization in an experimental setting are interesting research directions. Moreover, by interpreting the ring with a negative or fractional coefficient as a category, premodular fusion categories will appear, and they are expected to describe TOs in mixed states[273, 274, 275].